OpenStudy (anonymous):

i will fan and best anwser if you can calculate the problem on the first comment

OpenStudy (anonymous):

\[-1\frac{ 2 }{ 5 }\div \frac{ 5 }{ 8 }\]

OpenStudy (anonymous):

(-7/5) / (5/8)= -64/25

OpenStudy (ayovandera):

-0.64

OpenStudy (adamaero):

reported

OpenStudy (adamaero):

http://openstudy.com/code-of-conduct Don't devalue the question/answer process! Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem. Don't just provide the answer to a problem when someone else is in the middle of helping! But if you want to help, by all means, join in!

OpenStudy (anonymous):

wow adamaero

OpenStudy (anonymous):

@felavin dont you have to flip the last fraction then multiply?

OpenStudy (adamaero):